# Policy Gradient Optimzation for Bayesian-Risk MDPs with General Convex Losses

### v1

@@ 1 Introduction
Markov decision process (MDP) is a paradigm for modeling sequential decision making under uncertainty, with a primary focus on identifying an optimal policy that minimizes the (discounted) expected total cost. However, the standard form of MDP is not sufficient for modeling some practical problems. For example, consider a self-driving car operating in a dynamic urban environment. On one hand, the self-driving car must not only reach its destination efficiently but also safely against unpredictable events, requiring a general optimization objective within the MDP framework. On the other hand, the car has incomplete knowledge about its environment, such as road conditions. In such a case, the decision maker encounters two key challenges: the need for a general performance measure to address intrinsic uncertainty, and epistemic uncertainty about the environment. This paper is motivated by these challenges and aims to address both the general loss function and epistemic uncertainty simultaneously in the MDP framework.
@@
There is extensive literature addressing general loss functions and epistemic uncertainty separately. For instance, risk-sensitive objectives have been explored in the contexts of MDPs (Howard & Matheson, 1972; Ruszczyński, 2010; Mannor & Tsitsiklis, 2011; Petrik & Subramanian, 2012), stochastic optimal control (Borkar & Meyn, 2002; Moon, 2020), and stochastic programming (Shapiro, 2012; Pichler et al., 2022) literature. These objectives cannot be simply represented as the total expected cost. Epistemic uncertainty arises when some MDP parameters, such as transition probabilities, are unknown and must be estimated from available data. This discrepancy between the estimated and true MDP is referred to as epistemic uncertainty. Numerous approaches have been proposed to address epistemic uncertainty in MDPs, with robust MDP (Nilim & Ghaoui, 2004; Iyengar, 2005; Delage & Mannor, 2010; Wiesemann et al., 2013; Petrik & Russel, 2019) being one of the most widely adopted formulations. A more flexible and less conservative formulation, coined as Bayesian Risk MDP, was recently proposed by (Lin et al., 2022).
However, no existing literature addresses both a general loss function and epistemic uncertainty simultaneously. To the best of our knowledge, this paper is the first to consider this problem. In this work, we study MDPs with a general loss function, particularly focusing on convex functions in terms of the occupancy measure. Additionally, to handle both epistemic uncertainty and intrinsic uncertainty, we take a Bayesian approach to estimate the unknown parameters (such as transition probabilities) with a fixed batch of data, and impose a coherent risk functional (with respect to the Bayesian posterior distribution) on the general convex loss function. Therefore, the problem is framed as an offline optimization task with a two-layer composite objective: the outer general coherent risk measure and the inner general convex loss function. To determine the optimal policy for this composite problem, we use a policy optimization approach, which directly optimizes policies and accommodates complex, high-dimensional representations such as neural networks. This method typically employs parameterized policies and utilizes a policy gradient approach, introduced by (Sutton et al., 1999), to search for the optimal solution. For the outer layer, the coherent risk measure admits a dual representation as demonstrated by (Shapiro et al., 2021), which can be expressed as the supremum of the expectation over a risk envelope set. We extended the envelope theorem in (Milgrom & Segal, 2002) to obtain the policy gradient. A similar approach was taken by (Tamar et al., 2015), but their consideration of a discrete parameter space limits the applicability of their method to our problem. Our extension from the discrete case to the continuous case for the envelope theorem may be of independent interest. The derived policy gradient involves the gradient of the loss function with respect the policy parameter, which can then be estimated by different methods. In particular, we adapt the recent variational approach proposed by (Zhang et al., 2020) to construct the gradient estimator. Other methods, such as zeroth-order estimation method proposed by (Balasubramanian & Ghadimi, 2022), could also be used to estimate the inner gradient. By incorporating the inner gradient estimator into the policy gradient, we derive the gradient estimator for the composed objective and use policy gradient descent to optimize the problem. To make our approach more applicable with new observed data, we further extend our approach to the episodic setting, where the agent iteratively applies the current policy to gather more data and updates the policy based on new environment estimates informed by the additional data.
Our choice of policy gradient method for this problem is not only due to its popularity but also because algorithms based on dynamic programming are not applicable to general loss function that is not in the standard form of cumulative sum. Therefore, our approach is completely different from most robust MDPs or Bayesian risk MDPs which relies on dynamic programming. However, for feasibility of the policy gradient method, we assume the general loss function is convex. The convex loss functions are widely used, as discussed by (Pennings & Smidts, 2003), and are sufficiently general to encompass many of the previously mentioned examples (e.g., risk-sensitive MDPs and constrained MDPs) as special cases.More discussions about convex MDP is offered in Appendix A. The standard expected total cost can be viewed as such a special case, where the loss function is a linear function of the occupancy measure. The dynamic programming approach to solving MDPs involves the use of Bellman equations. However, the derivation of Bellman equations relies on the interchangeability principle, which may not hold for general convex loss functions. For a more detailed discussion on why the interchangeability principle fails for general convex loss functions, we refer readers to (Rockafellar & Wets, 2009) for the expectation operator and (Shapiro, 2017) for general risk functionals. It is also worth noting that the Bayesian approach has been considered by (Duff, 2002; Poupart et al., 2006; Abbasi-Yadkori & Szepesvári, 2015; Imani et al., 2018; Derman et al., 2020; Lin et al., 2022; Wang & Zhou, 2023), where the Bayesian update accounts for future data realization and enables the use of dynamic programming algorithms.
@@
For our composite problem, there have been dedicated efforts to solve MDPs with some special objectives using the policy gradient algorithm. For example, (Chow & Ghavamzadeh, 2014) applied Conditional Value-at-Risk(CVaR) to the total cost and developed policy gradient and actor-critic algorithms, each utilizing a distinct method to estimate the gradient and update policy parameters in the descent direction. In contrast, we consider a broader composition of a general coherent risk measure and a general loss function, allowing more flexible objectives. Note that although the composition of a coherent risk measure and a convex loss function is convex in the occupancy measure, it is generally non-convex in the policy parameters, which introduces additional challenges for our convergence analysis. Finally, the work most relevant to ours is perhaps (Zhang et al., 2020), which addresses a reinforcement learning (RL) problem with a general convex loss function and derives the variational policy gradient theorem with a global convergence guarantee. However, our work differs in that we consider an offline planning problem in an MDP with unknown transition probabilities, which are estimated from data. Therefore, we address not only a general convex loss function but also epistemic uncertainty. This introduces additional challenges related to risk measures, and the robustness of the proposed formulation is a key consideration.
@@
Our contributions are summarized as follows: (1) We propose a Bayesian risk formulation for MDPs with a general convex loss function and develop a policy gradient algorithm to solve for the optimal policy. The proposed formulation jointly mitigates both epistemic and intrinsic uncertainty; (2) We extend the envelope theorem to the dual representation of the coherent risk measure, and then apply the envelope theorem to derive the policy gradient. Our extension from the discrete case to the continuous case for the envelope theorem may be of independent interest; (3) We prove the convergence of the proposed algorithm and establish its convergence rate as $\mathcal{O}(T^{-1/2}+r^{-1/2})$, where $T$ is the number of policy gradient iterations and $r$ is the sample number of the gradient estimator; (4) We extend our policy gradient algorithm to the episodic setting, and prove the asymptotic convergence of the episodic minimizer of our Bayesian formulation to a global minimizer of the MDP problem under the true environment. Moreover, we show the number of iterations required in any episode to maintain an optimality gap $\mathcal{O}(\epsilon)$ under our Bayesian formulation.

### v2

# Bayesian Risk-Sensitive Policy Optimization For MDPs With General Loss Functions

@@ 1 Introduction
Markov decision process (MDP) is a paradigm for modeling sequential decision making under uncertainty, with a primary focus on identifying an optimal policy that minimizes the (discounted) expected total cost. However, the standard form of MDP is not sufficient for modeling some practical problems. For example, consider a self-driving car in a dynamic urban environment. It must reach its destination efficiently yet safely amid unpredictable events, calling for a general optimization objective within the MDP framework. At the same time, it faces incomplete knowledge of its environment, such as road conditions. Thus, the decision maker confronts two challenges: defining a general performance measure for intrinsic uncertainty and handling epistemic uncertainty about the environment. This paper addresses both challenges simultaneously within the MDP framework.
@@
There is extensive literature addressing general loss functions and epistemic uncertainty separately. For instance, risk-sensitive objectives have been explored in the contexts of MDPs (Howard & Matheson, 1972; Ruszczyński, 2010; Mannor & Tsitsiklis, 2011; Petrik & Subramanian, 2012), stochastic optimal control (Borkar & Meyn, 2002; Moon, 2020), and stochastic programming (Shapiro, 2012; Pichler et al., 2022) literature. These objectives cannot be simply represented as the total expected cost. Epistemic uncertainty arises when some MDP parameters, such as transition probabilities, are unknown and must be estimated from available data. This discrepancy between the estimated and true MDP is referred to as epistemic uncertainty. Numerous approaches have been proposed to address epistemic uncertainty in MDPs, with robust MDP (Nilim & Ghaoui, 2004; Iyengar, 2005; Delage & Mannor, 2010; Wiesemann et al., 2013; Petrik & Russel, 2019) being one of the most widely adopted formulations. A more flexible and less conservative formulation, coined as Bayesian Risk MDP, was recently proposed by Lin et al. (2022).
To the best of our knowledge, this paper is the first to address Markov decision processes (MDPs) with both general loss functions and epistemic uncertainty. Existing methods for handling epistemic uncertainty in MDPs (e.g., robust or Bayesian-risk formulations) typically rely on Bellman equations, which in turn depend on an interchangeability principle. This principle usually fails when the objective is a general loss function rather than the standard cumulative total cost, making dynamic programming inapplicable. Consequently, one cannot simply wrap a robust or Bayesian-risk method around a general objective.
Motivated by this limitation, we take a policy optimization approach to MDPs with general loss functions, particularly focusing on convex loss functions in the occupancy measure. Convexity is a natural assumption—such loss functions are widely used (see Pennings & Smidts (2003)) and encompass many existing models, such as risk-sensitive MDPs, constrained MDPs, and the standard expected total cost as a special case (where the loss is linear in the occupancy measure). To handle both epistemic uncertainty (unknown parameters like transition probabilities) and intrinsic uncertainty, we adopt a Bayesian perspective: we estimate parameters from a fixed batch of data and impose a coherent risk functional with respect to the Bayesian posterior distribution. This yields an offline optimization problem with a two-layer composite objective: an outer coherent risk measure and an inner convex loss function.
To solve this composite problem, we directly optimize policies and accommodate high-dimensional representations such as neural networks. The coherent risk measure admits a dual representation as shown in Shapiro et al. (2021), which can be expressed as a supremum over a risk envelope set. Building on this, we extend the envelope theorem in Milgrom & Segal (2002) to derive the policy gradient. Unlike Tamar et al. (2015), whose method was limited to a discrete parameter space, our extension covers the continuous case and may be of independent interest. The resulting policy gradient involves the gradient of the loss with respect to policy parameters, which we estimate using the recent variational method proposed in Zhang et al. (2020). Alternative estimators, such as the zeroth-order method of Balasubramanian & Ghadimi (2022), are also applicable. Incorporating these estimators into the policy gradient allows us to perform gradient descent on the composed objective. Finally, to enhance practicality, we extend our method to the episodic setting, where the agent iteratively gathers new data under the current policy and refines its estimates and policy updates accordingly.
@@
There have been efforts to solve some special cases of our composite problem using policy gradient algorithms. For example, Chow & Ghavamzadeh (2014) applied Conditional Value-at-Risk(CVaR) to the total cost and developed policy gradient and actor-critic algorithms, each utilizing a distinct method to estimate the gradient and update policy parameters in the descent direction. In contrast, we consider a broader composition of a general coherent risk measure and a general loss function, allowing more flexible objectives. Note that although the composition of a coherent risk measure and a convex loss function is convex in the occupancy measure, it is generally non-convex in the policy parameters, which introduces additional challenges for our convergence analysis. The work most relevant to ours is perhaps Zhang et al. (2020), which addresses a reinforcement learning (RL) problem with a general convex loss function and derives the variational policy gradient theorem with a global convergence guarantee. However, our work differs in that we consider an offline planning problem in an MDP with unknown transition probabilities, which are estimated from data. Therefore, we address not only a general convex loss function but also epistemic uncertainty. This introduces additional challenges related to risk measures, and the robustness of the proposed formulation is a key consideration.
@@
Our contributions are summarized as follows: (1) We propose a Bayesian risk formulation for MDPs with a general convex loss function and develop a policy gradient algorithm to solve for the optimal policy. The proposed formulation jointly mitigates both epistemic and intrinsic uncertainty; (2) We extend the envelope theorem to the dual representation of the coherent risk measure, and then apply the envelope theorem to derive the policy gradient. Our extension from the discrete case to the continuous case for the envelope theorem may be of independent interest; (3) We prove the convergence of the proposed algorithm and establish its convergence rate as $\mathcal{O}(T^{-1/2}+r^{-1/2})$, where $T$ is the number of policy gradient iterations and $r$ is the sample number of the gradient estimator; (4) We extend our policy gradient algorithm to the episodic setting, and prove the asymptotic convergence of the episodic optimal policy of our Bayesian formulation to a globally optimal policy of the MDP problem under the true environment. Moreover, we show the number of iterations required in any episode to maintain an optimality gap $\mathcal{O}(\epsilon)$ under our Bayesian formulation.